40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jun 23 21:35:46 2011, Jamie, Rana, Kiwamu, Update, SUS, Fix calibration for sus sensors 
    Reply  Thu Jun 23 22:00:22 2011, Jamie, Update, SUS, burt snapshot 
Message ID: 4868     Entry time: Thu Jun 23 21:35:46 2011     Reply to this: 4869
Author: Jamie, Rana, Kiwamu 
Type: Update 
Category: SUS 
Subject: Fix calibration for sus sensors 

We have fixed the counts-to-micron (cts2um) calibration for the suspension sensor filters. Each suspension sensor filter bank (e.g. ULSEN) has a "cts2um" calibration filter. These have now been set with the following flat gains:

   40 V       10^3 um         um
 -------- *  --------  = .36  --
 2^16 cts     1.7 V           ct

The INMTRX was also fixed with proper element values:

UL UR LR LL SIDE  
.25 .25 .25 .25 0 POS
1.666 1.666 -1.666

-1.666

0 PIT
1.666 -1.666 -1.666 1.666 0 YAW
0 0 0 0 1 SIDE

This was done for all core optic suspensions (BS, PRM, SRM, ITMX, ITMY, ETMX, ETMY).

 

ELOG V3.1.3-